OpenStudy (anonymous):

can someone check this? x+y=21 10.25x+13.50y=275.85

OpenStudy (anonymous):

so basically the question was asked how many white shirts were sold and how many yellow shirts were sold which was 2 white, and -19 yellow

Directrix (directrix):

Would you post the problem as it appears in your text or wherever? Thanks.

OpenStudy (anonymous):

okay sure

OpenStudy (anonymous):

one day a store sold 21 sweaters, white cost 10.95 and yellow 13.50 in all 275.85 shirts were sold how many of each sweater

Directrix (directrix):

> 275.85 shirts were sold Does this mean that $275.85 was the total amount of money made from the white and yellow shirts?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so i did 10.25(21-y)+13.50=275.85 first

OpenStudy (anonymous):

got 2125-10.25y+13.50y=275.85

Directrix (directrix):

Can someone check this? x+y=21--> ok 10.25x+13.50y=275.85 --> should be 10.95x + 13.50y = 275.85

OpenStudy (anonymous):

you are right. typo

Directrix (directrix):

10.95*(21-y)+13.50*y = 275.85

Directrix (directrix):

What did you get for y?

OpenStudy (anonymous):

229.95-10.95y+13.50=275.85?

Directrix (directrix):

You need to solve that equation for y.

Directrix (directrix):

This equation is not correct: 229.95-10.95y+13.50=275.85?

OpenStudy (anonymous):

im not sure what im doing wrong

Directrix (directrix):

Did you start with this equation: 10.95*(21-y)+13.50*y = 275.85

Directrix (directrix):

You have to get that typo out.

Directrix (directrix):

10.95* 21 - 10.95*y + 13.50*y = 275.85

OpenStudy (anonymous):

15

OpenStudy (anonymous):

15+6=21

Directrix (directrix):

229.95 + 2.55y = 275.85 2.55y = 275.85 - 229.95 2.55y = 45.9 y = ?

OpenStudy (anonymous):

18

Directrix (directrix):

y = 18 so the number of yellow shirts is 18. 21 - 18 = 3 so the number of white shirts is 3. Check the work.

OpenStudy (anonymous):

ok im walking through the steps again